OpenStudy (anonymous):

3. x over four + 10 = 14 (1 point) 96 46 16 –16 4. 3(4 – 2x) = –2x (1 point) –1 1 2 3

OpenStudy (anonymous):

A fraction means dividing the top number by the bottom number. So, put one of the answers in for x. 16 divided by 4 = 4 + 10 = 14!

OpenStudy (anonymous):

Thats not what its asking for...

OpenStudy (anonymous):

It is asking for x, isn't it? Then x = 16

OpenStudy (anonymous):

Yes. But not the way u did it.

OpenStudy (anonymous):

the simplified answer for #4 is 12 -6x = -2x

OpenStudy (anonymous):

I know the way you are talking about. But, that is the easy way I do it...

OpenStudy (anonymous):

Here is another way to do it: x/4+10 = 14 // - 14 x/4-14+10 = 0 1/4*x-4 = 0 // + 4 1/4*x = 4 // : 1/4 x = 4/1/4 x = 16 x = 16

OpenStudy (anonymous):

Ya ok that looks right haha sorry!! Thank u(:

OpenStudy (anonymous):

What about number 4?